OpenStudy (anonymous):

User: Casondra Bierman In Course: Algebra I V20 ( 4077) Instructor: Carlington Simms Exam: 04.06 Sequences Warning There is a checkbox at the bottom of the exam form that you MUST check prior to submitting this exam. Failure to do so may cause your work to be lost. Question 1 (Multiple Choice Worth 1 points) A store had 100 t-shirts. Each month, 30% of the t-shirts were sold and 25 new t-shirts arrived in shipments. Which recursive function best represents the number of t-shirts in the store, given that f(0) = 100?

OpenStudy (anonymous):

@BlazeRider2798 ????? any clue

OpenStudy (anonymous):

@tygrr321 im gonna suck you into this do you know?

OpenStudy (tygrr321):

?????????????????????????????????????????????????????

OpenStudy (anonymous):

f(n) = f(n − 1) ⋅ 0.3 + 25, n > 0 f(n) = 100 − f(n − 1) ⋅ 0.3 + 25, n > 0 f(n) = f(n − 1) ⋅ 0.7 + 25, n > 0 f(n) = 100 − f(n − 1) ⋅ 0.7 + 25, n > 0

OpenStudy (anonymous):

thats why i asked you

OpenStudy (anonymous):

@goldengoodgirl do you know?